# Embodied AI: Emerging Risks and Opportunities for Policy Action

### v1

@@ Abstract
The field of embodied AI (EAI) is rapidly advancing. Unlike virtual AI, EAI can exist in, learn from, reason about, and act in the physical world. Given recent innovations in large language and multimodal models, along with increasingly advanced and responsive hardware, EAI systems are rapidly growing in capabilities and operational domains. These advances present significant risks, including physical harm from malicious use, mass surveillance, and economic and societal disruption. However, these risks have been severely overlooked by policymakers. Existing policies, such as international standards for industrial robots or statutes governing autonomous vehicles, are insufficient to address the full range of concerns. While lawmakers are increasingly focused on AI, there is now an urgent need to extend and adapt existing frameworks to account for the unique risks of EAI. To help bridge this gap, this paper makes three contributions: first, we provide a foundational taxonomy of key physical, informational, economic, and social EAI risks. Secondly, we analyze policies in the US, EU, and UK to identify how existing frameworks address these risks and where these policies leave critical gaps. We conclude by offering concrete policy recommendations to address the coming wave of EAI innovation, including mandatory testing and certification for EAI systems, clarified liability frameworks, and forward-looking strategies to manage and prepare for transformative economic and societal impacts.
@@ 1 Introduction
EAI presents both opportunities and risks for humans. EAI systems already assist people with mobility impairments in navigating the world (e.g., autonomous cars), while future systems could fill crucial agricultural or manufacturing jobs as working-age populations decline. By augmenting and complementing human labor, EAI could foster significant economic development and prosperity Atkinson (2021). On the other hand, EAIs can more easily cause immediate physical damage than completely virtual AI systems and may cause significant social harm as humans and EAIs form closer connections (particularly with applications designed for companionship) Prescott and Robillard (2021); Oravec (2022). See Figure 1 for a schematic comparison of classical robots, agentic AI, and EAI.
Recent breakthroughs in AI capabilities—particularly those related to Large Language Models (LLMs) and Large Multimodal Models (LMMs)—have catalyzed unprecedented progress in EAIs’ ability to navigate and act in the physical world Kim et al. (2024); Gemini Robotics Team et al. (2025). At the same time, the rise of Vision-Language-Action Models (VLAs)—which cast control as next-token prediction over interleaved visual and linguistic tokens—opens the possibility for a “ChatGPT moment” for robotics, with sharp jumps in capability, deployment, and public awareness. Recent debuts of models like Gemini Robotics-ER, Alibaba’s Qwen2.5-VL, and NVIDIA’s Isaac GR00T N1 marked significant EAI algorithmic progress, even though these models are only slowly being paired with hardware advanced enough to translate virtual capabilities into real-world actions Gemini Robotics Team et al. (2025); Bai et al. (2025); noa (2025a). In the past few months, for example, EAIs have completed half-marathons and shown the ability to unpack groceries with little prior context Diviggiano and Baptista (2025); noa (2025b), and open-source resources from industry actors like Physical Intelligence and Unitree could spur continued technical progress noa (2025c); uni .
@@
EAI’s rapid growth in capabilities and deployment will increase the severity of potential harms and the urgency to address risks. This growth will necessitate significant updates to social, legal, and economic systems Suleyman and Bhaskar (2023); Floridi (2017). Although EAI shares many characteristics with virtual agentic AI Floridi and Sanders (2004), such as the varying degrees of autonomy and capability highlighted by Kasirzadeh et al. in “Characterizing AI Agents for Alignment and Governance” Kasirzadeh and Gabriel (2025), the physical embodiment of EAIs introduces distinct considerations and risks that warrant special attention Xing et al. (2025). EAIs can hit, cut, bump, maim, attack, and more, whether intentionally or unintentionally. The complexity of the physical world presents significant adaptation challenges for digital models trained in virtual simulations Semeraro et al. (2023); Mazumder et al. (2023). The coming wave of technological breakthroughs may also usher in a new era of scalability in which EAI will rapidly advance through an ever-improving software stack constrained by fewer human bottlenecks, thereby increasing the potential speed of change and compressing the timeline for action Wu et al. (2025).
@@ 2.1 Physical risks
Automation in sectors ranging from manufacturing to healthcare has and will increasingly put humans into close contact with EAI systems Garlick et al. (2024). This interaction increases the risk of accidental physical harm. Though accidental harm has been a longstanding issue in industrial robotics, increased AI capabilities could exacerbate this risk; several recent reports document an increase in industrial injuries following the introduction of AI-controlled robots Atkinson (2023); Fortune Staff (2023); Youn (2018). Virtual AI applications can also cause harm (e.g. through misinterpreted goals or misaligned behavior) but not directly in the physical world, unlike EAI. EAI’s potential to cause accidental, physical harm could be caused by misspecified goals, lack of semantic understanding, misaligned behavior, physical hardware malfunctions, or other unanticipated behaviors Amodei et al. (2016); Ji et al. (2025); Matheson et al. (2019). For example, a humanoid EAI might not correctly reason that placing a full glass of milk on a tilted table is perilous and likely to lead to a dangerous broken glass Fry , or a swarm of EAIs might get caught in a physical logjam and run over humans by mistake while trying to unstick themselves Winfield et al. (2025). Researchers also face persistent difficulty in getting models trained in purely virtual simulations to act as intended in the real physical world—what is referred to as the “reality gap.” Bousmalis et al. (2018). This introduces significant scope for accidental physical harm if the deployed world does not closely match an EAI’s training data Kaushik et al. (2022).
@@ 2.2 Informational risks
EAI systems interact with vast amounts of data, creating significant privacy concerns. EAIs are often trained on vast corpora and process a variety of data modalities—spanning visual, auditory, and tactile information—during deployment Kim et al. (2024). Like text-based virtual AI models, which are known to memorize and expose personally identifiable information Nasr et al. (2023); Carlini et al. (2019), commercial robots have been shown to disclose proprietary information through simple prompts Robey et al. (2025). Whereas virtual AI systems are constrained to collect data from either virtual interfaces or fixed points in the physical world (e.g. security cameras collecting facial-recognition data), EAI’s mobility and the vast array of sensors used in EAI technologies expand concerns about unauthorized data collection. For example, EAI systems can monitor user behavior, infer physical preferences, and potentially contribute to future model training without the informed consent of those being observed beyond the limitations of immobile microphones or security cameras Calo (2011); Chatzimichali et al. (2020); Guo (2022). Bad actors within governments or corporations could gain access to private data streams and monitor users’ movements 24/7, providing significant leverage over individuals to squash dissent or achieve personal power Davidson et al. (2025).
@@
Non-embodied AIs are known to propagate misinformation Chen and Shu (2024); Pan et al. (2023). Various studies have shown that LLMs hallucinate information, including academic citations Agrawal et al. (2023), clinical knowledge Singhal et al. (2023), and cultural references Weng (2024). EAIs inherit these shortcomings in the physical world, answering user questions with deceptive or incorrect information Danaher (2020). Because VLAs fuse vision and language, their hallucinatory failures can be spatially grounded—e.g., misidentifying an object in view and then generating a plausible yet unsafe action plan around it. And although automated home assistants like Amazon’s Alexa already lie about issues as innocuous as Santa Claus’ existence Withers (2018), more mobile, capable, and trusted EAIs in sensitive positions (like home-assistant or community-service positions) could easily spread model developers’ propaganda and talking points to users. For example, an EAI running on DeepSeek’s latest model could provide a subtle yet continuous stream of misinformation to American users while performing tasks as innocuous as folding laundry or helping to cook dinner Myers (2025); Lin (2024).
@@ 2.3 Economic risks
While virtual AI applications will likely displace certain types of human cognitive labor, EAIs could significantly replace or displace physical human labor Du et al. (2024). At a minimum, EAI will likely augment the type of work that humans perform Acemoglu and Autor (2011); Loaiza and Rigobon (2024). Classical industrial robots have taken over many human roles in manufacturing Grau et al. (2021), and research has shown that robot deployment can lead to a reduction in human employment Acemoglu and Restrepo (2020). Future technological advances will likely accelerate this displacement process, as increasingly capable EAIs perform complex, multi-step physical tasks beyond assembly lines—for instance, by serving as tourist guides or teaching in classrooms, and all without the need for sleep, breaks, sick leave, or vacation Newton and Newton (2019). Though automation has historically redirected labor toward areas of human comparative advantage Acemoglu and Restrepo (2019), AGI-enabled EAI could potentially automate all physical labor Korinek and Juelfs (2023).
@@
Along with displacing labor, EAI could significantly exacerbate wealth inequalities. Those who have access to or own EAI systems will be able to automate labor and perform many tasks significantly better or faster than those without access. These significant productivity advantages will potentially concentrate wealth and exacerbate domestic and international inequality Costinot and Werning (2018); Korinek and Stiglitz (2017). For example, while a wealthy businesswoman could invest in a fleet of the latest humanoid robots, individuals lacking adequate capital might be forced to rent their EAI systems. This division could create stark and entrenched socioeconomic divides as the importance of outsourcing labor to ever-more-capable EAIs increases Freeman (2016). Virtual AI applications may cause similar socioeconomic inequality, but the ability to use and control access to EAIs may confer significant and unique returns on investment, given that many tasks in the physical world necessary for human survival (e.g. growing food, building shelter) are constrained by human strength and energy.
@@
EAI deployment could accelerate the consolidation of economic and political power. Unlocking increasing returns to capital for EAI owners, EAI will decrease employers’ reliance on and responsiveness to the needs of human labor Kulveit et al. (2025). Further, EAI users or consumers may become dependent on EAI owners for goods and services due to the productivity advantages conferred by EAI systems Berg et al. (2018); Ford (2016). The importance of EAI to perform physical tasks will likely exacerbate power-concentration risks presented by purely virtual AI systems. The proliferation of EAIs could thus lead to a rapid concentration of corporate economic (and social) power, potentially even facilitating an eventual coup involving EAI Widder et al. (2024); Davidson et al. (2025).
@@ 2.4 Social risks
Like virtual applications of AI, EAIs can display bias towards and discriminate against users. When EAIs are placed in positions of power, their biases could have significant impacts on fairness in everyday interactions and on general social dynamics Howard and Borenstein (2018); Londoño et al. (2024). For example, a peacekeeping humanoid robot may discriminate based on skin color Azeem et al. (2024). Unlike virtual AI applications, this bias can have immediate and irreversible physical consequences (e.g. if the peacekeeping robot mistakenly injures an innocent passerby).
@@
Constant access to and interaction with EAIs could foster dangerous human dependence on or romantic attachment to EAIs Rabb et al. (2022). People may depend on EAIs for physical pleasure Cox-George and Bewley (2018). The physical presence and human-like features of EAI may significantly amplify the dependency issues already observed with conversational AI Fang et al. (2025); Freedman (2025). People may easily fall in love with EAIs, only to be distraught when these systems are altered or have their memories reset Jonze (2013).
@@
EAI deployment could fundamentally reshape society, particularly if the speed of technological development outpaces society’s ability to adapt Ford (2016); Brynjolfsson and McAfee (2014). For example, EAI systems could provide physical threats of violence and mass surveillance capabilities to back up AI-enabled authoritarianism Barez et al. (2025). Employers might only hire EAIs, leaving humans free to engage in other activities but also affecting how humans find meaning in their work Nikolova et al. (2024). Humans also might lose the ability to perform various tasks as responsibilities are increasingly outsourced or delegated to EAI systems Kulveit et al. (2025)—as an existing example, humans can lose natural abilities to navigate their environment when they outsource navigation to automated GPS systems Dahmani and Bohbot (2020). Of course, purely virtual AI may also transform human society (and, to some extent, already has). However, it is difficult to comprehend how profoundly and completely EAI systems could alter how humans work, socialize, and structure our societies—for example by revolutionizing physical labor systems, automating self-improvement and production, or becoming humans’ primary social and physical companions Boyd and Holton (2018).
@@ 3.1 Key policies
Beyond legislation, international standards provide manufacturers and deployers with robust safety guidance for robotics, primarily focused on physical safety. For example, ISO 10218:2025 recommends safety protocols for assessing risk, mitigating risk (e.g. through controls, safety functions, stopping functions), and safe-design certification for industrial robots noa (2025f). In addition, ISO 13482:2025 addresses safety requirements for service robots in personal and professional settings, emphasizing sensor reliability, uncertainty management protocols, and decision verification through multiple sensing modalities noa (2025g). Many standards also exist explicitly for AVs. For example, ISO/SAE 21434:2021 and UN Regulation 155 provide standards for cybersecurity engineering for road vehicles noa (2021a, b). These cybersecurity standards are particularly relevant for many forms of mobile EAIs, especially given that EAIs may be deployed in situations without regular access to trusted networks Thakker et al. (2025) (e.g. during disaster rescue missions, or when operating with law enforcement in hostile territory).
@@
GDPR legislation in the EU and UK establishes strict governance requirements concerning the capture, use, and storage of data. GDPR requires that data is only collected for “legitimate interests” and that entities collecting data are classified as “data controllers” who must document data collection practices, usage patterns, and storage methods while implementing robust security measures GDP (2016). These stipulations address many concerns with the privacy and treatment of data collected by EAIs. However, EAI deployments in public spaces challenge traditional consent models and controller identification Mintrom et al. (2022). Notions of data controllers and implied consent to be recorded will need to be substantially altered or clarified with EAI deployment. EAI systems deployed in public spaces, for example, raise questions about how to opt out of data collection, understand who receives and controls data collected from the system’s sensors, and even what constitutes a public vs. private space bar (2024).
@@
Few regulations directly address the social impacts of EAI systems. Those that do exist largely govern issues of direct human interaction with EAI and do not address larger issues of how society will transform as EAI systems become increasingly prevalent and powerful. The EU AI Act’s broad prohibition on infringing fundamental rights could be extended to address issues surrounding a lack of trust, lack of transparency, unhealthy or dangerous attachments, and bias and discrimination, but this would require further specification. In terms of accountability, proposed frameworks for attributing actions and delegating authority to virtual agents could prove helpful for EAIs Chan et al. (2025); South et al. (2025).
GDPR Article 22 provides an instructive example of existing regulation that implicates but does not directly address EAI systems. Likely designed with virtual AI applications in mind, the Article prohibits individuals from being “subject to a decision based solely on automated processing” when that decision has legal or similarly significant consequences for that individual GDP (2016). However, it remains unclear how this Article could be reconciled with fully autonomous EAIs, or how individuals could appeal to a human intervener—as the Article later mandates—in immediate physical interactions or conflicts with EAIs.

### v2

@@ Abstract
The field of embodied AI (EAI) is rapidly advancing. Unlike virtual AI, EAI systems can exist in, learn from, reason about, and act in the physical world. With recent advances in AI models and hardware, EAI systems are becoming increasingly capable across wider operational domains. While EAI systems can offer many benefits, they also pose significant risks, including physical harm from malicious use, mass surveillance, as well as economic and societal disruption. These risks require urgent attention from policymakers, as existing policies governing industrial robots and autonomous vehicles are insufficient to address the full range of concerns EAI systems present. To help address this issue, this paper makes three contributions. First, we provide a taxonomy of the physical, informational, economic, and social risks EAI systems pose. Second, we analyze policies in the US, EU, and UK to assess how existing frameworks address these risks and to identify critical gaps. We conclude by offering policy recommendations for the safe and beneficial deployment of EAI systems, such as mandatory testing and certification schemes, clarified liability frameworks, and strategies to manage EAI’s potentially transformative economic and societal impacts.
@@ 1 Introduction
EAI presents both opportunities and risks for humans. EAI systems already assist people with mobility impairments in navigating the world (e.g., autonomous cars), while future systems could fill crucial agricultural or manufacturing jobs as working-age populations decline. By augmenting and complementing human labor, EAI could foster significant economic development and prosperity Atkinson (2021). On the other hand, EAI systems can more easily cause immediate physical damage than completely virtual AI systems and may cause significant social harm as humans and EAI systems form closer connections (particularly with applications designed for companionship) Prescott and Robillard (2021); Oravec (2022). See Figure 1 for a schematic comparison of classical robots, agentic AI, and EAI.
Recent breakthroughs in AI capabilities—particularly those related to Large Language Models (LLMs) and Large Multimodal Models (LMMs)—have catalyzed unprecedented progress in EAI systems’ ability to navigate and act in the physical world Kim et al. (2024); Gemini Robotics Team et al. (2025). At the same time, the rise of Vision-Language-Action Models (VLAs)—which cast control as next-token prediction over interleaved visual and linguistic tokens—opens the possibility for a “ChatGPT moment” for robotics, with sharp jumps in capability, deployment, and public awareness. Recent debuts of models like Gemini Robotics-ER, Alibaba’s Qwen2.5-VL, and NVIDIA’s Isaac GR00T N1 marked significant EAI algorithmic progress, even though these models are only slowly being paired with hardware advanced enough to translate virtual capabilities into real-world actions Gemini Robotics Team et al. (2025); Bai et al. (2025); noa (2025a). In the past few months, for example, EAI systems have completed half-marathons and shown the ability to unpack groceries with little prior context Diviggiano and Baptista (2025); noa (2025b), and open-source resources from industry actors like Physical Intelligence and Unitree could spur continued technical progress noa (2025c); uni .
@@
EAI’s rapid growth in capabilities and deployment will increase the severity of potential harms and the urgency to address risks. This growth will necessitate significant updates to social, legal, and economic systems Suleyman and Bhaskar (2023); Floridi (2017). Although EAI shares many characteristics with virtual agentic AI Floridi and Sanders (2004), such as the varying degrees of autonomy and capability highlighted by Kasirzadeh et al. in “Characterizing AI Agents for Alignment and Governance,” Kasirzadeh and Gabriel (2025), the physical embodiment of EAI systems introduces distinct considerations and risks that warrant special attention Xing et al. (2025). EAI systems can hit, cut, bump, maim, attack, and more, whether intentionally or unintentionally. The complexity of the physical world presents significant adaptation challenges for digital models trained in virtual simulations Semeraro et al. (2023); Mazumder et al. (2023). The coming wave of technological breakthroughs may also usher in a new era of scalability in which EAI will rapidly advance through an ever-improving software stack constrained by fewer human bottlenecks, thereby increasing the potential speed of change and compressing the timeline for action Wu et al. (2025).
@@ 2.1 Physical risks
Automation in sectors ranging from manufacturing to healthcare has and will increasingly put humans into close contact with EAI systems Garlick et al. (2024). This interaction increases the risk of accidental physical harm. Though accidental harm has been a longstanding issue in industrial robotics, increased AI capabilities could exacerbate this risk; several recent reports document an increase in industrial injuries following the introduction of AI-controlled robots Atkinson (2023); Fortune Staff (2023); Youn (2018). Virtual AI applications can also cause harm (e.g. through misinterpreted goals or misaligned behavior) but not directly in the physical world, unlike EAI. EAI’s potential to cause accidental, physical harm could be caused by misspecified goals, lack of semantic understanding, misaligned behavior, physical hardware malfunctions, or other unanticipated behaviors Amodei et al. (2016); Ji et al. (2025); Matheson et al. (2019). For example, a humanoid EAI might not correctly reason that placing a full glass of milk on a tilted table is perilous and likely to lead to a dangerous broken glass Fry , or a swarm of EAI systems might get caught in a physical logjam and run over humans by mistake while trying to unstick themselves Winfield et al. (2025). Researchers also face persistent difficulty in getting models trained in purely virtual simulations to act as intended in the real physical world—what is referred to as the “reality gap.” Bousmalis et al. (2018). This introduces significant scope for accidental physical harm if the deployed world does not closely match an EAI’s training data Kaushik et al. (2022).
@@ 2.2 Informational risks
EAI systems interact with huge amounts of data, creating significant privacy concerns. These systems are often trained on vast corpora and process a variety of data modalities—spanning visual, auditory, and tactile information—during deployment Kim et al. (2024). Like text-based virtual AI models, which are known to memorize and expose personally identifiable information Nasr et al. (2023); Carlini et al. (2019), commercial robots have been shown to disclose proprietary information through simple prompts Robey et al. (2025). Whereas virtual AI systems are constrained to collect data from either virtual interfaces or fixed points in the physical world (e.g. security cameras collecting facial-recognition data), EAI’s mobility and the vast array of sensors used in EAI technologies expand concerns about unauthorized data collection. For example, EAI systems can monitor user behavior, infer physical preferences, and potentially contribute to future model training without the informed consent of those being observed beyond the limitations of immobile microphones or security cameras Calo (2011); Chatzimichali et al. (2020); Guo (2022). Bad actors within governments or corporations could gain access to private data streams and monitor users’ movements 24/7, providing significant leverage over individuals to squash dissent or achieve personal power Davidson et al. (2025).
@@
Non-embodied AIs are known to propagate misinformation Chen and Shu (2024); Pan et al. (2023). Various studies have shown that LLMs hallucinate information, including academic citations Agrawal et al. (2023), clinical knowledge Singhal et al. (2023), and cultural references Weng (2024). EAI systems inherit these shortcomings in the physical world, answering user questions with deceptive or incorrect information Danaher (2020). Because VLAs fuse vision and language, their hallucinatory failures can be spatially grounded—e.g., misidentifying an object in view and then generating a plausible yet unsafe action plan around it. And although automated home assistants like Amazon’s Alexa already lie about issues as innocuous as Santa Claus’ existence Withers (2018), more mobile, capable, and trusted EAI systems in sensitive positions (like home-assistant or community-service positions) could easily spread model developers’ propaganda and talking points to users. For example, an EAI running on DeepSeek’s latest model could provide a subtle yet continuous stream of misinformation to American users while performing tasks as innocuous as folding laundry or helping to cook dinner Myers (2025); Lin (2024).
@@ 2.3 Economic risks
While virtual AI applications will likely displace certain types of human cognitive labor, EAI systems could significantly replace or displace physical human labor Du et al. (2024). At a minimum, EAI will likely augment the type of work that humans perform Acemoglu and Autor (2011); Loaiza and Rigobon (2024). Classical industrial robots have taken over many human roles in manufacturing Grau et al. (2021), and research has shown that robot deployment can lead to a reduction in human employment Acemoglu and Restrepo (2020). Future technological advances will likely accelerate this displacement process, as increasingly capable EAI systems perform complex, multi-step physical tasks beyond assembly lines—for instance, by serving as tourist guides or teaching in classrooms, and all without the need for sleep, breaks, sick leave, or vacation Newton and Newton (2019). Though automation has historically redirected labor toward areas of human comparative advantage Acemoglu and Restrepo (2019), AGI-enabled EAI could potentially automate all physical labor Korinek and Juelfs (2023).
@@
Along with displacing labor, EAI could significantly exacerbate wealth inequalities. Those who have access to or own EAI systems will be able to automate labor and perform many tasks significantly better or faster than those without access. These significant productivity advantages will potentially concentrate wealth and exacerbate domestic and international inequality Costinot and Werning (2018); Korinek and Stiglitz (2017). For example, while a wealthy businesswoman could invest in a fleet of the latest humanoid robots, individuals lacking adequate capital might be forced to rent their EAI systems. This division could create stark and entrenched socioeconomic divides as the importance of outsourcing labor to ever-more-capable EAI increases Freeman (2016). Virtual AI applications may cause similar socioeconomic inequality, but the ability to use and control access to EAI systems may confer significant and unique returns on investment, given that many tasks in the physical world necessary for human survival (e.g. growing food, building shelter) are constrained by human strength and energy.
@@
EAI deployment could accelerate the consolidation of economic and political power. Unlocking increasing returns to capital for EAI owners, EAI will decrease employers’ reliance on and responsiveness to the needs of human labor Kulveit et al. (2025). Further, EAI users or consumers may become dependent on EAI owners for goods and services due to the productivity advantages conferred by EAI systems Berg et al. (2018); Ford (2016). The importance of EAI to perform physical tasks will likely exacerbate power-concentration risks presented by purely virtual AI systems. The proliferation of EAI systems could thus lead to a rapid concentration of corporate economic (and social) power, potentially even facilitating an eventual coup involving EAI Widder et al. (2024); Davidson et al. (2025).
@@ 2.4 Social risks
Like virtual applications of AI, EAI can display bias towards and discriminate against users. When EAI systems are placed in positions of power, their biases could have significant impacts on fairness in everyday interactions and on general social dynamics Howard and Borenstein (2018); Londoño et al. (2024). For example, a peacekeeping humanoid robot may discriminate based on skin color Azeem et al. (2024). Unlike virtual AI applications, this bias can have immediate and irreversible physical consequences (e.g. if the peacekeeping robot mistakenly injures an innocent passerby).
@@
Constant access to and interaction with EAI systems could foster dangerous human dependence or romantic attachment Rabb et al. (2022). People may depend on EAI systems for physical pleasure Cox-George and Bewley (2018). The physical presence and human-like features of EAI systems may significantly amplify the dependency issues already observed with conversational AI Fang et al. (2025); Freedman (2025). People may easily fall in love with EAI systems, only to be distraught when these systems are altered or have their memories reset Jonze (2013).
@@
EAI deployment could fundamentally reshape society, particularly if the speed of technological development outpaces society’s ability to adapt Ford (2016); Brynjolfsson and McAfee (2014). For example, EAI systems could provide physical threats of violence and mass surveillance capabilities to back up AI-enabled authoritarianism Barez et al. (2025). Businesses might only employ EAI systems, leaving humans free to engage in other activities but also affecting how humans find meaning in their work Nikolova et al. (2024). Humans also might lose the ability to perform various tasks as responsibilities are increasingly outsourced or delegated to EAI systems Kulveit et al. (2025)—as an existing example, humans can lose natural abilities to navigate their environment when they outsource navigation to automated GPS systems Dahmani and Bohbot (2020). Of course, purely virtual AI may also transform human society (and, to some extent, already has). However, it is difficult to comprehend how profoundly and completely EAI could alter how humans work, socialize, and structure our societies—for example by revolutionizing physical labor norms, automating self-improvement and production, or becoming humans’ primary social and physical companions Boyd and Holton (2018).
@@ 3.1 Key policies
Beyond legislation, international standards provide manufacturers and deployers with robust safety guidance for robotics, primarily focused on physical safety. For example, ISO 10218:2025 recommends safety protocols for assessing risk, mitigating risk (e.g. through controls, safety functions, stopping functions), and safe-design certification for industrial robots noa (2025f). In addition, ISO 13482:2025 addresses safety requirements for service robots in personal and professional settings, emphasizing sensor reliability, uncertainty management protocols, and decision verification through multiple sensing modalities noa (2025g). Many standards also exist explicitly for AVs. For example, ISO/SAE 21434:2021 and UN Regulation 155 provide standards for cybersecurity engineering for road vehicles noa (2021a, b). These cybersecurity standards are particularly relevant for many forms of mobile EAI, especially given that EAI systems may be deployed in situations without regular access to trusted networks Thakker et al. (2025) (e.g. during disaster rescue missions, or when operating with law enforcement in hostile territory).
@@
GDPR legislation in the EU and UK establishes strict governance requirements concerning the capture, use, and storage of data. GDPR requires that data is only collected for “legitimate interests” and that entities collecting data are classified as “data controllers” who must document data collection practices, usage patterns, and storage methods while implementing robust security measures GDP (2016). These stipulations address many concerns with the privacy and treatment of data collected by EAI systems. However, EAI deployments in public spaces challenge traditional consent models and controller identification Mintrom et al. (2022). Notions of data controllers and implied consent to be recorded will need to be substantially altered or clarified with EAI deployment. EAI systems deployed in public spaces, for example, raise questions about how to opt out of data collection, understand who receives and controls data collected from the system’s sensors, and even what constitutes a public vs. private space bar (2024).
@@
Few regulations directly address the social impacts of EAI. Those that do exist largely govern issues of direct human interaction with EAI systems and do not address larger issues of how society will transform as these entities become increasingly prevalent and powerful. The EU AI Act’s broad prohibition on infringing fundamental rights could be extended to address issues surrounding a lack of trust, lack of transparency, unhealthy or dangerous attachments, and bias and discrimination, but this would require further specification. In terms of accountability, proposed frameworks for attributing actions and delegating authority to virtual agents could prove helpful for EAI Chan et al. (2025); South et al. (2025).
GDPR Article 22 provides an instructive example of existing regulation that implicates but does not directly address EAI systems. Likely designed with virtual AI applications in mind, the Article prohibits individuals from being “subject to a decision based solely on automated processing” when that decision has legal or similarly significant consequences for that individual GDP (2016). However, it remains unclear how this Article could be reconciled with fully autonomous EAI, or how individuals could appeal to a human intervener—as the Article later mandates—in immediate physical interactions or conflicts with EAI systems.